OpenStudy (pgreene008):

Find the first six terms of the sequence. (1 point) a1 = -8, an = 5 • an-1

OpenStudy (pgreene008):

these are my options: -40, -200, -1000, -5000, -25,000, -125,000 0, 5, -40, -35, -30, -25 -8, -40, -200, -1000, -5000, -25,000 -8, -40, -35, -30, -25, -20

OpenStudy (jdoe0001):

any ideas what \(\Large a_{\color{brown}{ n}}=5\cdot a_{{\color{brown}{ n}}-1}\) means?

OpenStudy (pgreene008):

i mean thats what i have to use to plug in a1=8 right?

OpenStudy (jdoe0001):

hm?

OpenStudy (jdoe0001):

do you know hmmm what \(\Large a_{\color{brown}{ n}}=5\cdot a_{{\color{brown}{ n}}-1}\) that means by any chance?

OpenStudy (pgreene008):

sort of, not really. i was confused

OpenStudy (jdoe0001):

well... have you covered geometric sequences yet?

OpenStudy (pgreene008):

yeah

OpenStudy (jdoe0001):

alrite... so... lemme hmm hmmm post something

OpenStudy (jdoe0001):

\(\begin{array}{llllr} a_{\color{brown}{ n}}&=&5\cdot &a_{{\color{brown}{ n}}-1}\\ the\ {\color{brown}{ n}}^{th}\textit{ term}&equals&5\ times&previous\ term \end{array}\) does that make any sense? =)

OpenStudy (jdoe0001):

the notation \(\Large a_{\color{brown}{ n}}=5\cdot a_{{\color{brown}{ n}}-1}\) simply means, "the current term value, is 5 times (whatever the n-1 or this term minus one, or the previous one was)"

OpenStudy (jdoe0001):

which means that, to get the "next term", simply multiply your current term by 5 which means, 5 is the "common ratio" you're asked to find the next 6 terms well, you have the first one already, that is \(\Large a_1\) to get the others \(\begin{array}{llll} term&value \\\hline\\ a_1&-8\\ a_2&5\cdot a_1\\ a_3&5\cdot a_2\\ a_4&5\cdot a_3\\ a_5&5\cdot a_4\\ a_6&5\cdot a_4 \end{array}\)

OpenStudy (jdoe0001):

hmm \(\begin{array}{llll} term&value \\\hline\\ a_1&-8\\ a_2&5\cdot a_1\\ a_3&5\cdot a_2\\ a_4&5\cdot a_3\\ a_5&5\cdot a_4\\ a_6&5\cdot a_5 \end{array}\) rather